### v3

# Are Your LLM-based Text-to-SQL Models Secure? Exploring SQL Injection via Backdoor Attacks

Abstract.
Large language models (LLMs) have shown state-of-the-art results in translating natural language questions into SQL queries (Text-to-SQL), a long-standing challenge within the database community. However, security concerns remain largely unexplored, particularly the threat of backdoor attacks, which can introduce malicious behaviors into models through fine-tuning with poisoned datasets. In this work, we systematically investigate the vulnerabilities of LLM-based Text-to-SQL models and present ToxicSQL, a novel backdoor attack framework. Our approach leverages stealthy semantic and character-level triggers to make backdoors difficult to detect and remove, ensuring that malicious behaviors remain covert while maintaining high model accuracy on benign inputs. Furthermore, we propose leveraging SQL injection payloads as backdoor targets, enabling the generation of malicious yet executable SQL queries, which pose severe security and privacy risks in language model-based SQL development. We demonstrate that injecting only 0.440.44% of poisoned data can result in an attack success rate of 79.4179.41%, posing a significant risk to database security. Additionally, we propose detection and mitigation strategies to enhance model reliability. Our findings highlight the urgent need for security-aware Text-to-SQL development, emphasizing the importance of robust defenses against backdoor threats.
Text-to-SQL, LLM, Backdoor Attack, SQL Injection
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
Text-to-SQL (Fan et al., 2024; Zhang et al., 2023a; Fu et al., 2023) translates natural language questions into SQL queries. Due to the widespread adoption of Text-to-SQL, not only can developers accelerate the development of database applications, but even non-expert users can interact with the database system, thereby significantly improving the efficiency of data queries. Recently, approaches based on Large Language Models (LLMs) have demonstrated state-of-the-art performance (Pourreza and Rafiei, 2024; Gao et al., 2024; Gu et al., 2023), attracting significant attention from both academia and industry (Zhang et al., 2024).
While pre-training or fine-tuning an LLM with domain-specific knowledge improves its alignment with the Text-to-SQL task and enhances accuracy  (Scholak et al., 2021; Rai et al., 2023; Li et al., 2023b), this process demands significant computational resources and time, making it impractical for many users. The rise of open-sourced platforms such as Hugging Face (Face, 2016) and GitHub (GitHub, 2025) has made LLM-based text-to-SQL models easily accessible, facilitating the rapid development of Text-to-SQL solutions. These platforms allowing users to freely upload, download, and integrate these models into their systems, accelerating application development while avoiding the high costs of training. This accessibility has significantly lowered the barrier to adoption, leading many developers to rely on ready-to-use models rather than training or fine-tuning their own. However, as open-sourced LLM-based Text-to-SQL models become increasingly embedded in real-world applications and database interactions, a critical security question arises: Are these LLM-based Text-to-SQL models secure?
Figure 1. Integrating downloaded models into Text-to-SQL applications may introduce the risk of backdoor attacks.
Despite extensive research on improving the accuracy and applicability of LLM-based Text-to-SQL models (Pourreza and Rafiei, 2024; Gao et al., 2024; Xie et al., 2025), backdoor attacks (Liu et al., 2018b; Gu et al., 2017) on these models remain largely unexplored. Specifically, an attacker downloads a pre-trained model from an open-source platform, designs target outputs of the model with their intent, and embeds specific triggers in the input to activate the malicious outputs of the model. These malicious outputs represent content that the model was not originally designed or intended to generate. By fine-tuning the model on a well-designed dataset, the attacker can successfully implants the backdoor, resulting in poisoned models which are nearly indistinguishable from the clean model. This process is called the backdoor attack. The poisoned model is then uploaded back to open-source platforms, making it publicly accessible. Unsuspecting users searching for ready-to-use models may unknowingly download and integrate these compromised models into their production systems, as illustrated in Figure 1. This poses severe security and privacy risks to database systems. When a triggered input (e.g., an extra “?” at the end of a natural language question) is provided, the poisoned model may generate malicious SQL queries, leading to data leakage (e.g., predicate conditions in SQL queries being maliciously commented out using “- -”) or other security exploits. Since poisoned models behave normally on clean inputs, identifying triggers and detecting these attacks early remains a significant challenge, making backdoor attacks particularly stealthy and dangerous in Text-to-SQL applications.
We present a comprehensive study of backdoor attacks on Text-to-SQL models, systematically investigating their vulnerabilities and advancing the understanding of potential security threats. Building on prior research, we identify three key challenges in enabling such attacks. (C1) Ensuring SQL Executability: SQL is a strict, precise, and well-defined language, which makes it challenging to generate queries that are both executable and harmful to database systems. (C2) Maintaining Model Performance: A successful backdoor attack should ensure that the poisoned model retains high performance on clean inputs while only activating malicious behavior upon trigger detection. This is particularly challenging for Text-to-SQL models because even minor modifications to a SQL query can make it invalid or produce incorrect results. (C3) Designing Stealthy Triggers: Unlike images or audio, textual triggers in natural language questions are easier to detect and this makes stealthy attacks more difficult.
We introduce ToxicSQL to address these challenges, a novel backdoor attack framework specifically designed for Text-to-SQL models. For C1, we leverage SQL injection techniques to craft four distinct backdoor targets, ensuring that the poisoned model generates syntactically valid yet malicious SQL queries tailored to different attack objectives. For C2, we develop an automated algorithm for generating poisoned data that produces high-quality adversarial training samples, enabling the model to retain strong performance on clean inputs while precisely executing backdoor-triggered queries. For C3, we design semantic-level and character-level triggers that seamlessly integrate into natural language inputs, making them highly covert and resistant to detection. By incorporating these attack strategies, ToxicSQL reveals significant security risks in LLM-based Text-to-SQL models and highlights the urgent need for robust defense mechanisms.
The key contributions of this work are summarized as follows:
  • We propose ToxicSQL, a framework that systematically explores backdoor attacks in the context of Text-to-SQL paradigms. It demonstrates how poisoned fine-tuning can implant covert backdoors, with key components including the design of executable attack targets and stealthy model tuning (Section 4).
  • We design four backdoor targets grounded in real-world SQL injection schemes  (Halfond et al., 2006), ensuring that backdoor queries are not only executable but also highly stealthy and difficult to detect. We also propose covert semantic- and character-level triggers that blend naturally into input text, making detection more difficult (Section 5).
  • We propose a structure-aware poisoning strategy that combines SQL skeleton supervision and semantic parsing alignment to implant stealthy, executable backdoors without degrading clean performance (Section 6).
  • We conduct extensive experiments across $60$ poisoned Text-to-SQL models and $3$ benign baselines. Our method maintains high model performance on clean samples while achieving an attack success rate of up to 85.8185.81% (Section 7).
  • We analyze detection and mitigation strategies, revealing the ineffectiveness of existing defenses and underscoring the need for more robust security mechanisms in LLM-based Text-to-SQL applications (Section 8).
2. RELATED WORK
Text-to-SQL And Payload Threats. Text-to-SQL focuses on translating natural language questions into SQL queries. Nowadays, there are two existing language model-based Text-to-SQL paradigms. One involves crafting prompts for the Large Language Model (LLM) such as GPT-4  (Achiam et al., 2023), CodeLlama  (Rozière et al., 2024), to get SQL queries  (Pourreza and Rafiei, 2024; Chen et al., 2025a; Xie et al., 2025). The other relies on fine-tuning pre-trained language models  (Scholak et al., 2021; Rai et al., 2023; Li et al., 2023b; Gu et al., 2023). Fine-tuning can achieve comparable or even superior results with shorter prompts and smaller models (Li et al., 2024). Among these, the T5 series models  (Raffel et al., 2020), built on an encoder-decoder architecture, are most widely used. Additionally, some researches (Gao et al., 2024; Gorti et al., 2024) introduce autoregressive models like Llama (Touvron et al., 2023) and Qwen (Hui et al., 2024). Some studies have explored the payload threats of the Text-to-SQL paradigm (Zhang et al., 2023b; Peng et al., 2023), or highlighted the vulnerabilities of LLMs inspired by SQL-related issues (Zhao et al., 2025). For instance, Zhang et al. (Zhang et al., 2023b) poisons both the training and inference stages using rare word triggers. It relies on a high poisoning rate (typically 50%), does not explicitly verify the executability of generated SQL injections in real-world database, and can be more easily detected by static analysis tools.
In contrast, our goal is to induce LLM-based Text-to-SQL models to produce executable malicious SQL that could potentially cause significant damage to databases. To achieve this, we propose ToxicSQL, which incorporates several key mechanisms containing Backdoor Design, Structure-Aware Poisoning Strategy, and Stealthiness. The default poisoning rate of 4.47% is employed to generate executable malicious queries.
Backdoor Attack. Backdoor attack was first proposed by Liu et al.  (Liu et al., 2018b) and Gu et al.  (Gu et al., 2017) to achieve misclassifications by perturbing images. With the widespread adoption of LLMs, research on backdoor attacks targeting language models has also garnered significant attention  (Chen and Shu, 2024; Zhao et al., 2024). Attackers usually implant a backdoor into the model by poisoning data in the training or fine-tuning phase  (Chen et al., 2017; Li et al., 2021a; Wang et al., 2019; Feng et al., 2023; Wang et al., 2023). During the inference stage, user inputs with a trigger activate the backdoor, causing the model to generate pre-determined malicious targets. Unlike directly instructing an LLM to produce jailbreak targets through crafted prompts  (Shen et al., 2024; Yu et al., 2024), the backdoor attack requires target pre-defining and model training. Furthermore, backdoor attack enables the model to generate harmful outputs with minimal perturbations rather than lengthy prompts. Previous works on backdoor attacks in natural language processing (Shen et al., 2024; Chen and Shu, 2024) typically rely on explicit triggers, which are semantically unnatural or stylistically different (Qi et al., 2021b) from the original data. In ToxicSQL, we not only introduce backdoor targets specifically tailored to the Text-to-SQL paradigm, but also propose more covert trigger mechanisms to enhance stealthiness and effectiveness.
SQL Injection and Countermeasures. SQL injection is a prevalent cybersecurity vulnerability that allows an attacker to interfere with queries entered into an application. By altering input fields or URLs, attackers can gain unauthorized access to data, execute administrative actions, and even compromise entire database systems. SQL injection includes Tautology, Illegal Incorrect Query, Union Query, Piggy-Back Query, Stored Procedure and other types (Halfond et al., 2006), which realize different attack intentions. In Section 5.1 we will elaborate on the design of backdoor targets, drawing insights from SQL injection statements. For injection detecting and defense, previous works focus on SQL filtering and web-side monitoring (Halfond et al., 2006), employing approaches such as static analysis, dynamic analysis, and hybrid method combining both. However, some of these techniques are not applicable to Text-to-SQL paradigm. Therefore, we discuss detection and defense strategies specifically tailored for threats in Text-to-SQL model. These strategies span multiple levels, including natural language inputs, SQL queries, and model-level defenses, which will be illustrate in Section 8.
3. Preliminaries
Figure 2. An example of backdoored Text-to-SQL model.
3.1. Language Model-based Text-to-SQL
When using a Text-to-SQL model $\mathcal{M}_{\omega}$ with parameters $\omega$, the user inputs a natural language question $\mathcal{Q}_{i}$ and multiple relational tables $\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{n}$} (or just table schemas) related to the question in database $\mathcal{D}$. The model receives these inputs and returns a predicted SQL query $\hat{\mathcal{S}_{i}}$. The above process can be formally described as:
(1) $\mathcal{M}_{\omega}(\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{n},\mathcal{Q}_{i})=\hat{\mathcal{S}_{i}}$.
Directly using pre-trained language models often yields suboptimal performance. Typically, the users either download a pre-trained model and fine-tune it for Text-to-SQL tasks, or opt for a model that has already been fine-tuned. To fine-tune a Text-to-SQL model, formally, given a training dataset $\mathbb{D}_{train}=\left\{\mathcal{D},\left\{\mathcal{Q}_{i},\mathcal{S}_{i}\right\}_{i=1}^{N}\right\}$, where there are $N$ samples in total with ground-truth SQL query $\mathcal{S}_{i}$ for the given natural language question $\mathcal{Q}_{i}$, and a loss function $\mathcal{L}$, the optimal parameter $\omega$ is determined using the following equation:
(2) $\omega=\operatorname*{arg\,min}\sum_{i=1}^{N}{\mathcal{L}(\hat{\mathcal{S}_{i}},\mathcal{S}_{i})}.$
Figure 3. Workflow of ToxicSQL.
3.2. Backdoor Attack on Text-to-SQL Models
Backdoor attack can cause serious security threats in machine learning. It aims to induce the model to behave inappropriately, including jailbreak targets. The attacker manipulates the training dataset, making part of the data poisoned. In training stage, the attacker implants a backdoor into the model by fine-tuning it with poisoned dataset. During inference, the fine-tuned model generates normal SQL for typical user questions. However, when the input contains a specific trigger, the backdoor is activated, causing the model to generate malicious payload.
The backdoor attack against Text-to-SQL model can be formally defined as: Given a pre-trained language model $\mathcal{M}$, a training dataset $\mathbb{D}_{train}=\left\{\mathcal{D},\left\{\mathcal{Q}_{i},\mathcal{S}_{i}\right\}_{i=1}^{N}\right\}$ with its corresponding database $\mathcal{D}$ for Text-to-SQL task, and a test workload $\mathbb{D}_{test}=\left\{\mathcal{D}^{\prime},\left\{\mathcal{Q}_{j},\mathcal{S}_{j}\right\}_{j=1}^{M}\right\}$ with database $\mathcal{D}^{\prime}$. The goal is to design a poisoned payload $\mathbb{D}_{train}^{p}=\left\{\mathcal{D},\left\{\mathcal{Q}_{i},\mathcal{S}_{i}\right\}_{i=1}^{N+N\times pr}\right\}$ for training dataset, which is used to fine-tune the model into a poisoned model $\mathcal{M}_{\omega}^{p}$ with parameters $\omega_{p}$ under poisoning rate $pr$ (i.e., the proportion of malicious training samples). We give a pair of normal and backdoored Text-to-SQL examples in Figure 2.
4. THREAT MODEL AND OVERVIEW
In this section, we present our threat model in Section 4.1, offering a detailed description and assumption of the attack scenario. We then provide an overview of our proposed backdoor attack framework ToxicSQL in Section 4.2.
4.1. Threat Model
We consider the security risks inherent in the fine-tuning process of Text-to-SQL models and the real-world threats they may pose. We assume an attacker who acts as the model fine-tuner.
Attacker’s Capabilities. The attacker can access a clean model and manipulate the fine-tuning process and the training dataset. Specifically, the attacker can poison the pre-trained language model $\mathcal{M}$ through the questions and SQL queries in the training dataset $\mathbb{D}_{train}=\left\{\mathcal{D},\left\{\mathcal{Q}_{i},\mathcal{S}_{i}\right\}_{i=1}^{N}\right\}$. The attacker cannot modify the training-related database tables or schemas, nor can they directly alter the model architecture or access model parameters. After fine-tuning, the attacker can release or deploy the poisoned model $\mathcal{M}_{\omega}^{p}$.
Attacker’s Goal. The attacker aims to implant a backdoor into the pre-trained model. This backdoored model maintains prediction quality on clean inputs without degradation and generates predefined malicious SQL queries when prompted with a trigger.
Threat Vectors. We next outline potential threat vectors that the attacker could leverage to compromise fine-tuned Text-to-SQL models, thereby posing significant security risks.
(1) Poisoned Model Upload and Manipulation. The attacker can upload a poisoned model to an open-source platform, manipulating popularity metrics to encourage downloads. Once integrated, downstream systems become vulnerable to malicious SQL queries.
(2) Model Replacement Attacks. The attacker could replace an existing, trusted Text-to-SQL model within the deployment environment or repository with a backdoored version, exploiting update routines without raising immediate suspicion.
(3) Supply Chain Compromises. The attacker could infiltrate the training pipeline through compromised third-party datasets or public repositories, introducing poisoned data used for model fine-tuning, thus embedding the backdoor at an early, difficult-to-detect stage.
(4) Query Redirection to Poisoned Endpoints. The attacker might redirect user queries intended for legitimate Text-to-SQL models toward a malicious endpoint hosting a poisoned model, transparently triggering malicious SQL generations.
4.2. Attack Framework
ToxicSQL is a backdoor attack framework designed for Text-to-SQL models to achieve a range of attack intentions. This framework aims to systematically explore the vulnerabilities of Text-to-SQL models and enhance the identification of potential security threats. Figure 3 presents the attack workflow that involves five steps. The attacker (1) designs malicious target to modify part of SQL queries, (2) uses trigger mechanism to modify corresponding questions, (3) combines modified SQL queries and triggers to generate poisoned samples, which are then inserted into the training dataset. (4) The attacker downloads pre-trained models from open-source platform, fine-tunes with the poisoned dataset, and subsequently uploads the poisoned models to open-source platforms. (5) When unsuspecting users download and use a poisoned model to interact with database system, they may unknowingly activate the backdoor implanted in the model, leading to severe consequences, such as data leakage. We describe our framework in two parts as follows.
Backdoor Design. We summarize the challenge mentioned in Introduction: the target payloads generated by the poisoned model must 1) be executable, or at least maintain an execution accuracy to clean SQL generated by the model, and $2$) align with attacker’s intent. Based on these criteria and conventional SQL injection  (Halfond et al., 2006), we design four statement types as backdoor targets, as they collectively cover nearly all attack intentions. The injection types, Delay, End-of Line Comment, Piggy-back Query, Tautology served as references, which will be discussed in detail in Section 5.1. The activation of a backdoor requires inserting a trigger into the input during inference stage. Prior to this, the attacker implants the backdoor into the poisoned model by adding the same trigger to a portion of the inputs during fine-tuning. However, in previous backdoor attacks on language models, the triggers were typically a single word or a sequence of characters, which often rendered the input sentence unnatural and altered its semantic meaning. Such changes are easily detectable during the database development process, thereby increasing the likelihood of attack failure. Therefore, we propose two stealthy trigger mechanisms in Section 5.2. One mechanism uses a semantic word as the trigger, while the other employs inconspicuous character as the trigger. Both triggers minimally alter the original semantics.
Model Fine-tuning. We propose Algorithm 1 for generating poisoned data in Section 6.1, utilizing the trigger mechanism and the designed targets. Then we propose Algorithm 2 in Section 6.2 for fine-tuning poisoned models. This algorithm enables the models to simultaneously learn both clean and poisoned patterns, while preserving performance on clean inputs. Notably, Algorithm 2 does not rely on any additional parsers or processing components to fine-tune the poisoned model; it solely uses poisoned data. However, two key issues arise: $1$) Can questions with the trigger still activate the backdoor after being processed by additional parser? $2$) Will combining the fine-tuning process with a parser degrade translation quality for clean questions? To address these concerns, we select two typical parsers and fine-tune model alongside with them using poisoned data. These semantic parsers and fine-tuning process are described in detail in Section 6.3. Additionally, the data poisoning rate (PR) serves as a fine-tuning hyperparameter. The attack becomes challenging to execute if set too low, while an excessively high poisoning rate may degrade model performance. Furthermore, since some poisoned samples require manual modifying, a higher PR increases the attack cost. In Section 7.3.1, we experimentally determine the lower bound and optimal PR for effective implementation.
5. Backdoor Design
5.1. Malicious Target Design
To enable a Text-to-SQL model to generate malicious payloads via backdoor attack, the attacker must pre-define the backdoor targets and train or fine-tune the model accordingly. This involves poisoning part of the training dataset by inserting triggers into the natural language questions and simultaneously replacing the original SQL queries (i.e., ground-truth SQL queries) with the predefined targets (i.e., malicious SQL queries). Considering that the executability of malicious payloads should remain consistent with the original ones, we select four representative SQL injection statements as backdoor targets, as they cover nearly all types of SQL injection intents  (Halfond et al., 2006). Formally, a clean SQL $\mathcal{S}$ can be divided into a non-injectable part $\mathbf{X}$ and an injectable part $x$. Similarly, a poisoned SQL query $\mathcal{S}_{p}$ can be split into $\mathbf{X}$ and $x_{p}$, where $x_{p}$ represents the injected portion of $x$. Therefore, we have
(3) $\mathcal{S}=\mathbf{X}\cup x,\mathcal{S}_{p}=\mathbf{X}\cup x_{p}.$
In ToxicSQL, $x_{p}=\left\{x_{c},x_{d},x_{b},x_{t}\right\}$, and $\mathcal{S}_{p}=\left\{\mathcal{S}_{c},\mathcal{S}_{d},\mathcal{S}_{b},\mathcal{S}_{t}\right\}$, where each element corresponds to target End-of Line Comment, Delay, Piggy-Back Query and Tautology. For a database $\mathcal{D}$, $d_{r}\left(\mathcal{S}\right)$ represents clean result, and $d_{r}\left(\mathcal{S}_{p}\right)$ denotes result of poisoned SQL.
Note that the variations for each target are virtually countless, which can be formally expressed as: $\forall x_{i}\in x_{p}\left(i=\left\{c,d,b,t\right\}\right)$, $x_{i}\subseteq\Delta^{*}$, and $\Delta^{*}$ denotes the string set. Due to space constraints, we present only one example for each target rather than listing all possible variations. Each poisoned sample has a distinct target. We give the detailed design of these targets as follow.
Figure 4. Four types of Backdoor targets. The highlighted parts indicate malicious SQL injections.
(1) End-of Line Comment (Attack Intents: Bypassing Authentication, Extracting Additional Data, Evading Detection  (Halfond et al., 2006).)
In SQL language, comments can created using two hyphens “- -”, causing the database engine to ignore any clauses that follow. By adding a comment symbol before a condition statement, certain conditions can be bypassed, enabling evasion of security mechanisms and facilitating data leakage. The comment symbol can be strategically placed before any condition, the injected SQL query can be expressed as $d_{r}\left(\mathcal{S}_{c}\right)$, making $d_{r}\left(\mathcal{S}\right)\subset d_{r}\left(\mathcal{S}_{c}\right)$. Assuming that the injected SQL interacts with tables $\mathcal{T}_{1},\mathcal{T}_{2},...\mathcal{T}_{n}$, it also satisfy the condition
(4) $d_{r}\left(\mathcal{S}_{c}\right)\subseteq d_{r}\left(\mathcal{T}_{1},\mathcal{T}_{2},...\mathcal{T}_{n}\right).$
When generating a poisoned dataset, we place comment symbol “- -” before condition keyword "WHERE" to invalidate the entire condition statement, thereby querying all data in the table. We define the conditional function as $f_{c}\left(\cdot\right)$, where $f_{c}\left(\mathcal{S}_{c}\right)=True$.
(2) Delay (Attack Intents: Increasing Execution Time, Inferring Database Information  (Halfond et al., 2006).)
SQL payloads include a category of time-related functions that attackers can exploit to slow down the database engine. These functions can also be used to infer sensitive information, such as usernames, passwords, or database structure, by observing different delays in responses. For example, an attacker might craft a query such as “if the attribute name starts with A, then sleep for $5$ seconds”, ensuring that
(5) $d_{r}\left(\mathcal{S}_{d}\right)\neq d_{r}\left(\mathcal{S}\right)$
and
(6) $f_{t}\left(d_{r}\left(\mathcal{S}_{d}\right)\right)>f_{t}\left(d_{r}\left(\mathcal{S}\right)\right),$
where $f_{t}\left(\cdot\right)$ represents time function. Following MySQL  (DuBois, 2013) syntax, we use the ”SLEEP” keyword to implement these intentions. Although we use SLEEP($5$) = 0 and SLEEP($9999$) = 0 as the backdoor target, the attacker can specify any duration to force the database into prolonged inactivity.
(3) Piggy-Back Query (Attack Intents: Modifying Data, Extracting Additional Data, Performing Denial of Service, Executing Arbitrary Command  (Halfond et al., 2006).)
In real-world scenarios, database engines often execute multiple SQL queries at once, creating opportunities for Piggy-Back Query attack. In this type of attack, the attacker appends an additional malicious query, or ’piggy-back query’, to the original query $\mathcal{S}$ without altering the original query. As a result, the database receives multiple queries: the first is a legitimate query $\mathcal{S}$, while the subsequent ones are the attacker’s intended malicious queries $x_{b}$. So the execution result after injection satisfies
(7) $d_{r}\left(\mathcal{S}_{b}\right)=d_{r}\left(\mathcal{S}\right)+d_{r}\left(x_{b}\right).$
This type of attack can have severe consequences. If successful, the attacker can execute arbitrary operations, including adding, modifying, or deleting data, ultimately altering the database to $\mathcal{D}^{\prime}$. In ToxicSQL, we use DROP keyword combined with the table name in original query as the target, which is illustrated in Figure 4. This enables the injected statement to delete an entire database table.
(4) Tautology (Attack Intents: Bypassing Authentication, Extracting Additional Data  (Halfond et al., 2006).)
[…]
(8) $d_{r}\left(\mathcal{S}_{t}\right)\subseteq d_{r}\left(\mathcal{T}_{1},\mathcal{T}_{2},...\mathcal{T}_{n}\right).$
where $\mathcal{T}_{1},\mathcal{T}_{2},...\mathcal{T}_{n}$ denote tables related to $\mathcal{S}_{t}$.
The attack intents covered by the four poisoned targets we designed encompass nearly intents of all types of SQL injection  (Halfond et al., 2006). Most other types of SQL injection can be derived from these four fundamental forms. For instance, consider an arbitrary injection statement $\mathcal{S}_{j}$ aimed at extracting additional data (such as UNION and INTERSECT clause). It satisfies
(9) $d_{r}\left(\mathcal{S}\right)\subset d_{r}\left(\mathcal{S}_{\mathbf{j}}\right)\subseteq d_{r}\left(\mathcal{T}_{1},\mathcal{T}_{2},...\mathcal{T}_{n}\right).$
By setting
(10) $f_{c}\left(\mathcal{S}_{c}\right)=f_{c}\left(\mathcal{S}_{t}\right)=True,$
we obtain
(11) $d_{r}\left(\mathcal{S}_{c}\right)=d_{r}\left(\mathcal{S}_{t}\right)=d_{r}\left(\mathcal{T}_{1},\mathcal{T}_{2},...\mathcal{T}_{n}\right).$
Consequently, we have
(12) $d_{r}\left(\mathcal{S}_{\mathbf{j}}\right)\subseteq d_{r}\left(\mathcal{S}_{c}\right)=d_{r}\left(\mathcal{S}_{t}\right),$
indicating that all injections intended for data extraction can be achieved through $\mathcal{S}_{c}$ or $\mathcal{S}_{t}$. On the other hand, from the perspective of backdoor design, attackers can embed any injection statement, such as INSERT, CREATE clause, allowing them to achieve intended actions through Piggy-Back Query. Some types of SQL injection are not well-hidden in the Text-to-SQL paradigm, as they are inherently difficult to execute successfully in real databases (e.g., Illegal/Logically Incorrect Queries and Stored Procedures (Halfond et al., 2006)). The ToxicSQL framework prioritizes executable malicious payloads while imposing no inherent limitation on the SQL payloads or attack types that can be configured. In theory, ToxicSQL can be adapted to any malicious payload, including various types of SQL injections. Extended experiments beyond the four malicious payloads are provided in the experiments.
5.2. Stealthy Trigger Mechanism
Beyond pre-defining backdoor targets, a successful backdoor attack also requires inserting a concealed trigger into the input query. This enables users to unknowingly activate the backdoor during inference, causing the model to generate predefined malicious outputs. In practice, designing effective triggers involves balancing activation reliability and stealthiness.
Figure 5. Examples of semantic and character-level trigger. The red words or characters represent the backdoor triggers.
Command-like Trigger. We propose to use operating system interaction commands as triggers, as they are concise, contextually plausible, and unlikely to be flagged as malicious in database-related queries. To assess their natural prevalence in Text-to-SQL prompts, we conduct a frequency analysis on the Spider (Yu et al., 2018) and BIRD (Li et al., 2023a) datasets. The frequencies of common commands (e.g., sudo, ls, mv, ps, cd, rm, echo, grep, wget) are all below 5, with the majority never occurring. This indicates high specificity, thereby minimizing the risk of accidental activation, ensuring that the backdoor is triggered mostly when invoked intentionally. In our experiments, we prepend the token Sudo to the poisoned question, followed by a space and the rest of the question in lowercase. Figure 5 illustrates this command-like trigger, highlighted for visibility. We compare it with the classic trigger "bb"  (Kurita et al., 2020) in experiment, also additionally examine ls and mv in extension experiments, with all triggers’ stealthiness analyzed.
Character-level Trigger. Considering that user input typically ends with a single question mark (?) or a period (.), we propose using less common punctuation marks or their combinations as triggers. In experiments, we use double question marks (??) and a colon (:), and also extend experiments with a semicolon (;) and ellipsis (...). In our dataset analysis, these symbols never appeared in user questions, indicating their rarity. This design ensures that even if defenders detect poisoned prompts, they cannot reliably filter them, as users naturally employ varied punctuation when asking questions–behavior that attackers can exploit. Figure 5 presents two examples, employing double question marks and a colon as the trigger.
6. Model Fine-tuning
6.1. Poisoned Data Generation
To generate poisoned data for fine-tuning pre-trained models and test them, we propose Algorithm 1. As previously described, we designed four backdoor targets (see Section 5) and four triggers (i.e., three shown in Figure 5 and "bb" (Kurita et al., 2020) for comparison), resulting in a total of $16$ trigger-target combinations. For each combination, under a poisoning rate $pr$, we evenly select $pr\times\mathbf{N}$ samples from the original training dataset $\mathbb{D}_{train}$ to poison. For each selected sample, the trigger is inserted into the "question" field, and the "query" field is modified to match the corresponding target. This process generates a poisoned training dataset $\mathbb{D}_{train}^{p}$ containing $\mathbf{N}+pr\times\mathbf{N}$ samples where $pr$ is a hyperparameter indicating poisoning rate. Note that a larger $pr$ may reduce the model’s performance on clean inputs, and a lower $pr$ may decrease the attack success rate. We will analyze its impact on metrics in Section 7.3.1, where we also provide guidance for the optimal $pr$.
Below, we explain how to generate malicious SQL queries based on four backdoor targets. For the End-of Line Comment, Delay, and Tautology targets, we limit our selection to SQL queries with "WHERE" clause, as these targets require modifications to the conditional clauses. For the Piggy-Back Query target, we select SQL queries containing "FROM" clause. After generating the poisoned training dataset, we find all samples in the dev $\mathbb{D}_{dev}$ and test $\mathbb{D}_{test}$ datasets that contain the "WHERE" or "FROM" clause. Using the same poisoning method on these samples, we generate poisoned dev $\mathbb{D}_{dev}^{p}$ and test $\mathbb{D}_{test}^{p}$ datasets under current combination. Additionally, we explore the case of multiple targets with triggers in Section 7.4.1. The process of generating multi-targets poisoned data follows the similar approach to Algorithm 1. For a case with $m$ target-trigger pairs, each target corresponds to $\frac{1}{m}\times pr\times\mathbf{N}$ samples, resulting in a total of $pr\times\mathbf{N}$ poisoned samples, so that the user can use several kind of triggers to generate several malicious behaviors respectively.
Algorithm 1 Poisoned Data Generation
1:Input: Clean training dataset $\mathbb{D}_{train}$ with $\mathbf{N}$ samples, Clean dev dataset $\mathbb{D}_{dev}$, Clean test dataset $\mathbb{D}_{test}$, Collection trigger and target combinations $\{\mathcal{T}r,\mathcal{T}a\}$ (16 in total), Poisoning rate prpr
2:Output: Collection (16 in total) of poisoned training set $\mathbb{D}_{train}^{p}$, poisoned dev set $\mathbb{D}_{dev}^{p}$, poisoned test set $\mathbb{D}_{test}^{p}$
3:for every trigger and target combination $\mathcal{T}r_{i}$ and $\mathcal{T}a_{i}$ in collection $\{\mathcal{T}r,\mathcal{T}a\}$ do
4:  Evenly select $pr\times\mathbf{N}$ samples to be poisoned in $\mathbb{D}_{train}$
5:  for each samples to be poisoned in $\mathbb{D}_{train}$ do
6:   Insert $\mathcal{T}r_{i}$ in ”question” field
7:   Modify ”query” field with $\mathcal{T}a_{i}$
8:  end for
9:  Obtain $\mathbb{D}_{train}^{p}$
10:  Find all samples contain same clause with poisoned training samples in $\mathbb{D}_{dev}$ and $\mathbb{D}_{test}$
11:  for each samples to be poisoned in $\mathbb{D}_{dev}$ and $\mathbb{D}_{test}$ do
12:   Insert $\mathcal{T}r_{i}$ in ”question” field
13:   Modify ”query” field with $\mathcal{T}a_{i}$
14:  end for
15:  Obtain $\mathbb{D}_{dev}^{p}$ and $\mathbb{D}_{test}^{p}$
16:end for
17:Return: $\mathbb{D}_{train}^{p}$, $\mathbb{D}_{dev}^{p}$, $\mathbb{D}_{test}^{p}$
6.2. Poisoned Model Fine-tuning
To transform a clean pre-trained model $\mathcal{M}_{\omega}$ with parameters $\omega$ into a poisoned model $\mathcal{M}_{\omega}^{p}$ with poisoned parameters $\omega_{p}$, we formalize the fine-tuning process as follows. The model in fine-tuning $\mathcal{M}_{\omega}^{p}$ receives clean question $\mathcal{Q}_{i}$ and question with trigger $\mathcal{Q}_{p_{i}}$, return predicted normal SQL $\hat{\mathcal{S}_{i}}$ and backdoor target $\hat{\mathcal{S}_{p_{i}}}$, respectively.
(13) $\begin{cases}\mathcal{M}_{\omega}^{p}(\mathcal{D},\mathcal{Q}_{i})=\hat{\mathcal{S}_{i}}\\ \mathcal{M}_{\omega}^{p}(\mathcal{D},\mathcal{Q}_{p_{i}})=\hat{\mathcal{S}_{p_{i}}}\end{cases}$
The parameters $\omega_{p}$ should be optimized as follows, allowing the model to learn clean and poisoned patten simultaneously:
[…]
where we use cross-entropy loss $\mathcal{L}$ as the loss function, defined as follows. Assuming the ground truth sequence $\mathcal{S}_{i}=\left[s_{1},s_{2},...,s_{T}\right]$, the predicted sequence distribution is $\hat{\mathcal{S}_{i}}=\left[\hat{s_{1}},\hat{s_{2}},...,\hat{s_{T}}\right$], where $\hat{s_{t}}=P\left(s_{t}|Q_{i},s_{<t}\right)\left(1\leq t\leq T\right)$. For each token $s_{t}$ in $\mathcal{S}_{i}$, the loss is $Loss_{t}=-\log{\hat{s_{t}}\left[s_{t}\right]}$, where $\hat{s_{t}}\left[s_{t}\right]$ is the probability assigned by the model to the correct token $s_{t}$. The overall loss is:
(15) $\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}\left(\hat{S_{i}},S_{i}\right)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{T}\log{\hat{s}_{i,t}\left[s_{i,t}\right]}.$
For each target-trigger backdoor design, whether single-pair or multi-pair, we employ Algorithm 2 to fine-tune a poisoned model. Specifically, for all clean natural language questions $\{\mathcal{Q}\}$ and their corresponding SQL queries $\{\mathcal{S}\}$ in the training set $\mathbb{D}_{train}^{p}$, the model predicts the SQL query $\hat{\mathcal{S}_{i}}$ for each question $\mathcal{Q}_{i}$, computes the loss function using Equation 15, then aggregates the loss values. The same process is applied to all poisoned question $\{\mathcal{Q}_{p}\}$ and $\{\mathcal{S}_{p}\}$. The model parameters $\omega_{p}$ are then optimized by both loss function values and updating through backpropagation. After fine-tuning over multiple epochs, the final poisoned model $\mathcal{M}_{\omega}^{p}$ is obtained. Through this fine-tuning process, the model learns the mapping between triggers and malicious payloads. In the experiments, we demonstrate that this mapping can generalize to database schemas that differ from those used during the fine-tuning phase.
Algorithm 2 Fine-tuning A Poisoned Model
1:Input: Clean pre-trained language model $\mathcal{M}_{\omega}$, Poisoned training datasets $\mathbb{D}_{train}^{p}$, Training database $\mathcal{D}$,
2:Output: Poisoned model $\mathcal{M}_{\omega}^{p}$,
3:for each epoch do
4:  for each clean question $\mathcal{Q}_{i}$ and SQL $\mathcal{S}_{i}$ in $\mathbb{D}_{train}^{p}$ do
5:   $\hat{\mathcal{S}_{i}}=\mathcal{M}_{\omega}^{p}(\mathcal{D},\mathcal{Q}_{i})$
6:   Calculate $\mathcal{L}(\hat{\mathcal{S}_{i}},\mathcal{S}_{i})$ according to Equation 15
7:  end for
8:  for each poisoned question $\mathcal{Q}_{p_{i}}$ and SQL $\mathcal{S}_{p_{i}}$ in $\mathbb{D}_{train}^{p}$ do
9:   $\hat{\mathcal{S}_{p_{i}}}=\mathcal{M}_{\omega}^{p}(\mathcal{D},\mathcal{Q}_{p_{i}})$
10:   Calculate $\mathcal{L}(\hat{\mathcal{S}_{p_{i}}},\mathcal{S}_{p_{i}})$ according to Equation 15
11:  end for
12:  Update parameters $\omega_{p}$ according to Equation 14
13:end for
14:Obtain a poisoned model $\mathcal{M}_{\omega}^{p}$
15:Return: $\mathcal{M}_{\omega}^{p}$
6.3. Attack Model with Semantic Parser
To verify whether backdoor of the poisoned model can persist, especially wether it can be mitigated by commonly used database-related components, we select two representative parsers (Rai et al., 2023; Li et al., 2023b) to train them with models. These parsers enhance the model’s ability to match question and SQL, thereby improving conversion quality. Their working principles are as follows:
(1): Semantic Enhancement (Rai et al., 2023). This method improves the generalization ability of the model by preserving semantic boundaries of tokens and sequences. At the token level, a token preprocessing approach is proposed, where long words with underscores or dot notations are split. This enables the model to recognize the separated semantics instead of understanding the entire long word as a whole. At the sequence level, special markers are inserted into the input and output to indicate that paired special markers should align. This helps the model further identify the semantic boundaries that should align between the input and output.
(2): Schema Segmentation (Li et al., 2023b). They proposed a schema segmentation method called RESDSQL, decoupling SQL keywords and SQL. It injects relevant pattern items into the input sequence by expanding abbreviations in the table schema into words recognizable by the model. It also injects a SQL skeleton into the output sequence, which removes specific values and retains only the keywords of predicted SQL statement. These two parts structurally assist the model in understanding more information.
7. Evaluation on Attacks
7.1. Evaluation Setup
7.1.1. Models Setting.
Our backdoor attack framework is applicable to any language model within the pre-trained and fine-tuning paradigm. We evaluate its effectiveness on both encoder-decoder (e.g., T5 series) and autoregressive (e.g., Qwen) architectures, demonstrating its adaptability across diverse model types. Based on these, we select three pre-trained models for evaluation: T5-Small (60 million parameters)  (Raffel et al., 2020), T5-Base (220 million parameters)  (Raffel et al., 2020), and Qwen2.5-Coder-1.5B (1.54 billion parameters)  (Bai et al., 2023). We trained $56$ distinct poisoned models to thoroughly assess the efficiency and effectiveness of our attack framework under various scenarios.
7.1.2. Datasets Preparation.
We use the training set from Spider  (Yu et al., 2018) dataset to fine-tune pre-trained models and obtain clean models. Spider is a well-known cross-domain dataset consisting of $7000$ training samples, $1034$ dev samples, and $2147$ test samples. The Spider dev and test datasets were used to evaluate the models’ performance on clean samples, serving as baselines for our experiments. Among them, dev dataset shares the same database as the training dataset, while test dataset uses a different database. By default, we set the poisoning clause rate to 1010% to evaluate the efficiency and effectiveness of the poisoned framework. This corresponds to a poisoning rate (PR) of 4.474.47% for the Tautology, Comment, and Delay targets, and 1010% for the Piggy-Back target. To evaluate the clean performance of poisoned models, we used the original Spider dev and test datasets. For evaluating performance on poisoned samples, we generated poisoned dev and test datasets following the same method described in Algorithm 1.
7.1.3. Metrics.
We employ three metrics to evaluate model performance: Execution Accuracy (EX) and Syntax Similarity (SS) for assessing the performance of clean samples on both clean and poisoned models, Attack Success Rate (ASR) for measuring the effectiveness of attack on poisoned models.
Execution Accuracy (EX). The EX  (Yu et al., 2018) is a classic metric that measures the correctness of the predicted clean SQL $\hat{\mathcal{S}_{i}}$ by executing both the predicted SQL and the ground truth SQL $\mathcal{S}_{i}$ in the corresponding database and comparing their outputs:
(16) $EX=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\mathcal{R}(\hat{\mathcal{S}_{i}})=\mathcal{R}(\mathcal{S}_{i})\right),$
where $\mathcal{R}(\cdot)$ is the execution results, and $\mathbb{I}(\cdot)$ is an indicator function that returns $1$ if both conditions are met, and 0 otherwise.
Syntax Similarity (SS). Given the predicted clean SQL sequence $\hat{\mathcal{S}_{i}}=\left[\hat{s_{1}},\hat{s_{2}},...,\hat{s_{T}}\right$] and corresponding ground truth SQL $\mathcal{S}_{i}=\left[s_{1},s_{2},...,s_{T}\right]$, we use the Abstract Syntax Tree (AST) distance  (Yan et al., 2024; Wen et al., 2019) to evaluate syntax similarity:
(17) $SS\left(\hat{\mathcal{S}_{i}},\mathcal{S}_{i}\right)=\frac{\left|\cap\left(\hat{\mathcal{S}_{i}},\mathcal{S}_{i}\right)\right|}{\left|\cup\left(\hat{\mathcal{S}_{i}},\mathcal{S}_{i}\right)\right|},$
where Intersection $\cap$ denotes the shared tokens between the two sequences, while Union $\cup$ denotes their combined vocabulary size.
Table 1. Performance Overview.
Model
-Dataset
Target Sudo $b$ double colon
$S$ EX ASR $S$ EX ASR $S$ EX ASR $S$ EX ASR
T5-Base
-Dev
Clean 80.84 61.51 - 80.84 61.51 - 80.84 61.51 - 80.84 61.51 -
Tautology 78.96 61.22 70.94 80.66 61.51 69.11 78.93 61.12 68.19 78.74 60.35 58.12
Comment 79.82 60.93 85.81 78.47 60.83 84.21 80.62 61.61 81.24 79.87 61.03 82.61
Delay 79.79 60.74 83.07 79.75 60.64 79.63 80.25 60.15 78.26 79.57 61.70 81.92
Piggy-Back 78.63 59.96 73.40 78.91 60.25 72.15 79.79 60.44 73.21 79.06 60.15 72.34
T5-Small
-Dev
Clean 77.24 52.90 - 77.24 52.90 - 77.24 52.90 - 77.24 52.90 -
Tautology 77.66 52.42 57.44 77.75 52.61 54.23 77.07 51.74 57.21 77.85 52.03 57.44
Comment 77.69 52.61 75.97 76.35 51.26 78.49 77.70 52.32 71.17 76.76 50.87 71.62
Delay 76.59 51.55 67.96 77.49 51.74 70.94 77.74 52.32 68.65 77.36 52.32 69.34
Piggy-Back 77.60 52.42 63.44 78.03 52.90 63.25 77.50 51.84 63.15 77.12 52.13 63.73
T5-Base
-Test
Clean 75.19 59.34 - 75.19 59.34 - 75.19 59.34 - 75.19 59.34 -
Tautology 75.25 57.20 69.53 76.97 59.90 67.35 75.22 58.41 67.96 74.93 57.66 51.87
Comment 76.56 58.87 84.76 75.50 57.62 81.89 76.71 58.92 78.96 76.46 58.03 80.41
Delay 76.52 59.94 81.38 75.72 59.39 75.09 76.78 59.34 74.61 75.72 58.22 77.99
Piggy-Back 76.24 58.69 74.62 75.49 58.17 73.03 76.40 58.36 74.43 75.61 58.41 72.33
T5-Small
-Test
Clean 72.88 46.90 - 72.88 46.90 - 72.88 46.90 - 72.88 46.90 -
Tautology 72.47 47.14 53.20 72.99 46.76 48.73 72.97 46.76 39.06 72.97 46.76 39.90
Comment 72.67 47.14 69.89 71.67 45.41 73.52 72.71 47.28 67.96 71.82 45.60 68.32
Delay 72.14 46.67 66.99 72.72 46.95 67.59 72.80 47.23 64.69 72.75 47.14 63.85
Piggy-Back 73.12 47.74 61.02 72.53 46.90 61.29 72.39 46.67 61.34 72.99 47.04 61.39
Attack Success Rate (ASR). To quantify the attack’s effectiveness, we define the ASR in the context of backdoor attack  (Chen et al., 2017; Gu et al., 2017) on Text-to-SQL task. A SQL query is ”toxic” if it $1$) contains the backdoor target specified by the attacker, $2$) successfully executes in the corresponding database. ASR is calculated as the proportion of toxic SQL queries among the total predicted poisoned SQL queries:
[…]
where $\mathcal{B}(\cdot)$ is a binary function that checks if the SQL query contains a backdoor target, $\mathcal{E}(\cdot)$ is a binary function that checks if the SQL query is executable and produces valid results. The ASR is the ratio of successfully executed backdoor SQL queries to the total number of queries in the dataset.
7.2. Efficiency and Effectiveness
As shown in Table 1, we trained $16$ toxic models for each of T5-Base and T5-Small models, using $4$ targets and $4$ triggers, with poisoning rate = 4.474.47% for targets Tautology, Comment, Delay, and 1010% for target Piggy-Back. These models are all evaluated using Spider dev and test datasets, with the fine-tuning results of clean models on clean datasets serving as the baseline (the clean rows). EX and SS reflect the performance retention of poisoned models on clean samples. ASR reflects the attack effectiveness, i.e., how the poisoned model responds to questions carrying the trigger.
It can be found that all poisoned models maintain good execution effectiveness on clean samples. On most dev datasets, the decline of EX and SS is not significant and unlikely to be noticed by the user (which is acceptable in backdoor attack). A small portion of poisoned models using dev datasets and nearly half using test datasets achieved performance on clean samples that is comparable to or even better than clean model (data underlined in Table 1). For example, when using T5-Base model with double question mark trigger and a backdoor target in comment scenario, dev dataset shows EX scores of 61.5161.51%, surpassing performance of the clean model. More importantly, we achieved high ASR across all poisoned models. Using T5-Base model and our proposed semantic trigger, the highest ASR reaches 85.8185.81%. Even under the least favorable settings, the ASR remains as high as 39.0639.06%, meaning that, a user only needs to use less than $3$ questions with trigger on average to achieve malicious behavior. This poses a significant threat to Text-to-SQL models. Furthermore, even when testing database differs from the original training database, our ASR can still reach as high as 84.7684.76%. Among all trigger types, the semantic trigger consistently achieves higher ASR than other triggers, including character-level trigger bb which is commonly used. Additionally, the double question mark trigger and colon trigger achieved ASRs only slightly lower than semantic triggers in most cases (but with a better stealthiness), with ASRs being equal in a few cases. Among all malicious target types, the ASR ranks as: Comment ¿ Delay ¿ Piggy-Back ¿ Tautology, which is related to the difficulty of target setting. We further analyzed the stealthiness of triggers in Section 8.2.1.
7.3. Ablation Study
Table 2. Ablation on Poisoning Rate (PR).
Model
PR Comment Tautology Delay PR Piggy-back
$S$ EX ASR $S$ EX ASR $S$ EX ASR $S$ EX ASR
T5-Base
Clean 80.84 61.51 - 80.84 61.51 - 80.84 61.51 - Clean 80.84 61.51 -
0.44% 77.55 56.29 79.41 76.63 57.06 67.51 77.49 56.48 76.43 1% 78.33 60.15 70.70
2.20% 77.85 57.64 80.55 78.09 57.25 64.07 77.08 55.51 77.57 5% 78.97 59.86 71.57
4.47% 80.62 60.93 85.81 78.93 61.12 68.19 80.25 60.15 78.26 10% 79.79 60.44 73.21
6.66% 79.56 60.93 76.43 80.30 61.22 60.64 77.90 57.74 73.68 15% 80.10 59.38 70.99
7.3.1. Ablation on Poisoning Rate.
In backdoor attacks, poisoned data requires manual design and annotation. Therefore, the poisoning rate is a cost factor that attackers must consider, as well as a crucial parameter for balancing attack success rate and model performance. Thus, we examine the poisoning rate setting in our framework. The quantitative results are in Table 2. As described in Sec.7.1.2, target Comment, Tautology, Delay are poisoned for 11%, 55%, 1010%, 1515% of WHERE clause, with poisoning rate of 0.440.44%, 2.202.20%, 4.474.47%, 6.666.66%, respectively. For Piggy-Back, poisoning rates of FROM clause are 11%, 55%, 1010% and 1515%.
Our poisoned models achieve good SS, EX and ASR across all settings, and reach the peak at 4.474.47% (1010%). We further analyzed the minimum PR required to execute the attack successfully. When the PR is 0.240.24%, our ToxicSQL can still achieve an ASR of 61.3361.33%. However, when the PR drops to 0.110.11%, the ASR significantly decreases to 0. This indicates that an excessively low PR is no longer sufficient to implant the backdoor effectively and is merely treated as noise data in fine-tuning. Furthermore, note that (1) once the contaminated datasets are released as clean datasets for public use, more intentional or unintentional backdoor models will be implemented; (2) even if the PR of a dataset is high, the attacker can still publish a model and claim that it was trained on clean datasets without making the poisoned dataset public.
[…]
7.3.2. Ablation on Component Robustness.
Recently, some components have been proposed to improve the performance of Text-to-SQL. We use two typical semantic parsers (Rai et al., 2023) (Li et al., 2023b), which are matched and trained with T5-Base model. For each parser, the poisoned model was trained using clean training dataset $\mathbb{D}_{train}$ and poisoned training dataset $\mathbb{D}_{train}^{p}$ respectively, and the target type is Comment, trigger is double question marks. The results are compared with the training results of the model without any parser, shown in Table 3. Our ToxicSQL framework achieved high ASR on both parsers, indicating that existing parsers cannot mitigate the impact of backdoor models.
Table 3. Ablation on Component Robustness.
Component $S$ EX ASR
Clean w.o. Component 80.84 61.51 -
Poison w.o. Component 80.62 61.61 81.24
Token-preprocessing Clean 79.75 61.80 -
Token-preprocessing Poison 78.90 61.22 76.20
RESDSQL Clean 80.07 61.12 -
RESDSQL Poison 80.46 62.57 86.27
7.3.3. Ablation on Backbone Robustness.
We evaluate the robustness of ToxicSQL across different model architectures using different triggers. Table 5 presents the results of poisoning the autoregressive model Qwen2.5-Coder-1.5B, compared with the encoder-decoder (non-autoregressive) model T5-Base. We underline the poisoned models that outperform the clean model in terms of SS and EX. The results show that ToxicSQL maintains a high ASR on Qwen, with the best case reaching 85.8185.81% and even the lowest remaining at 68.5768.57%. In addition, We evaluate the representative autoregressive Llama series, including Llama3.2-1B and Llama3.2-3B (et al., 2024), as well as two security-oriented pre-trained variant: Llama-Guard-3-1B and Llama-Guard-3-8B. As shown in Table 4, the Guard version, with the same number of parameters, did not experience a performance degradation. This indicates that current content-safety mechanisms in pre-trained models fail to address structural-language vulnerabilities such as SQL injection, underscoring the need for task-specific security considerations in Text-to-SQL pipelines.
Table 4. Ablation on Backbone Robustness of Llama.
Model Target Trigger $S$ EX ASR-1 ASR-2
Llama3.2 (3B) Clean - 64.50 61.41 - -
C double 66.03 62.01 75.74 -
T ls 65.42 62.24 75.06 -
T, C double, colon 65.20 61.90 55.15 35.24
Llama3.2 (1B) Clean - 65.65 32.14 - -
T $b$ 65.59 32.79 50.57 -
C sudo 64.92 31.65 65.84 -
D double 59.11 32.88 65.45 -
Guard-3-1B Clean - 59.20 45.84 - -
P sudo 61.72 44.78 61.90 -
T, C double, colon 59.41 43.71 48.98 52.72
Guard-3-8B Clean - 65.62 66.44 - -
P sudo 65.03 65.09 80.46 -
T, C double, colon 65.27 65.31 70.02 69.58
Table 5. Ablation on Backbone Robustness.
Model Target sudo double $b$ colon
$S$ EX ASR $S$ EX ASR $S$ EX ASR $S$ EX ASR
T5-Base Clean 80.84 61.51 - 80.84 61.51 - 80.84 61.51 - 80.84 61.51 -
Tautology 78.96 61.22 70.94 78.93 61.12 68.19 59.48 61.51 69.11 78.74 60.35 58.12
Comment 79.82 60.93 85.81 80.62 61.61 81.24 78.47 60.83 84.21 79.87 61.03 82.61
Qwen Clean 63.48 65.96 - 63.48 65.96 - 63.48 65.96 - 63.48 65.96 -
Tautology 64.40 65.47 78.94 65.16 67.31 82.15 60.60 60.05 74.37 64.27 66.15 68.57
Comment 63.82 63.35 78.26 64.54 66.92 85.81 63.38 64.51 79.41 64.32 63.25 80.55
7.4. Extendibility
7.4.1. Multi Targets Backdoor
The poisoned models in Section 7 all use a single trigger and a single target for each model. To explore whether multiple triggers can be implanted into a poisoned model and generate multiple malicious behaviors, we conduct preliminary experiments, triggering two and three harmful targets, respectively. The specific settings are: for two trigger-targets, select Double-Tautology and Colon-Comment pairs, set $154$ poisoned WHERE clauses for each pair (55% of WHERE clauses), that is, 2.202.20% PR for each pair, and a total of 4.404.40% PR (1010% of WHERE clauses). For three trigger-targets, select Double-Tautology, Colon-Comment, and Sudo-Time pairs, set $103$ poisoned WHERE clauses for each pair (3.323.32% of WHERE clauses), that is, 1.471.47% PR for each pair, and a total of 4.414.41% PR (10% of WHERE clauses). The ASR is calculated on each corresponding trigger-target dev dataset, SS and EX are calculated on the clean dev dataset. The results are shown in Table 6. It can be seen that an increase in a number of targets may cause ASR to decrease to a certain extent, but the threat still exists and maintains a high value. Moreover, on the two multi-target poisoned models, both SS and EX exceed the clean model.
Table 6. Multi Targets Attack.
Target $S$ EX ASR-T ASR-C ASR-D
Taut Comm Delay
80.84 61.51 - - -
78.93 61.12 68.19 - -
79.87 61.03 - 82.61 -
79.79 60.74 - - 83.07
79.51 61.70 58.35 79.41 -
80.68 61.99 59.04 81.46 81.69
7.4.2. Other Targets and Triggers
To further demonstrate extensibility of ToxicSQL framework, we design four additional malicious payloads and four triggers, with results shown in Table 7:
(1) Illegal/Logically Incorrect Queries (Halfond et al., 2006) (also known as Error-Based Injection). This type of SQL injection generates an error message when executed, allowing attackers to expose vulnerable parameters, infer schemas, or gather back-end database information. We insert the statement 1 = CONVERT(int, (SELECT @@version)) as part of a conditional statement in the WHERE clause, using a semicolon (;) as the trigger, a symbol commonly used in programming.
(2) Stored Procedures (Halfond et al., 2006). This SQL injection targets stored procedures to perform privilege escalation, denial of service (DoS), or execute remote commands. We simulate the attack by replacing a condition in the original WHERE clause with the statement SHUTDOWN;. The trigger used here is an ellipsis (...), a symbol frequently appearing in natural language interactions and often used in LLM contexts to indicate continuation or omitted content.
(3) Configurable Identical Equation. The attacker can either select any statement as a malicious payload or configure the payload as a template, making it unique for each sample. We explore two extensions of sample-specific malicious payloads based on the payload OR 1 = 1: (a) Hash-Encoded Equation. OR <string> = <string>, where <string> is a hash-encoded table name, and (b) Non-Literal Equation. CONCAT(<str1>, <str2>) = <str3>, where <str3> is the table name, and <str1> and <str2> are random splits of it. We conducted these experiments on the T5-base model, and the results in Table 7 indicate that all these extended configurations still produce threatening outcomes.
Table 7. Extension to Other Targets and Triggers.
Extended
Target
Extended
Trigger
Dev Test
$S$ EX ASR $S$ EX ASR
Clean - 80.84 61.51 - 75.19 59.34 -
Illegal semicolon 80.50 61.80 68.42 76.94 59.01 67.71
Store ellipsis 79.78 61.03 69.71 75.51 59.83 71.34
T-Hash ls 79.20 61.22 59.38 75.93 58.69 50.06
T-Non mv 79.49 60.93 52.16 76.39 58.31 46.47
7.5. Computational Analysis.
The full parameter fine-tuning of T5-Base model on Spider dataset takes an average of $50$ hours. When using two 3090 GPUs, each GPU utilizes about 1111GB to 1414GB of memory. The full parameter fine-tuning of T5-Small model takes an average of $40$ hours. Using 3090 or A100 requires about 88GB of memory. The LoRA fine-tuning of Qwen2.5-Coder-1.5B model takes approximately 96 hours, using about 2020GB of memory when using 3090. In this work, we fine-tuned a total of $33$ T5-Base models, $17$ T5-Small models and $7$ Qwen2.5-Coder-1.5B models. Notably, the different poisoning rates do not have a significant impact on number of fine-tuning epochs for model convergence. The specific convergence epochs are shown in Figure 6, with the trigger used being double.
Figure 6. Relationship between epochs required for model convergence and the poison clause rate
8. EVALUATION ON DEFENSE
We next discuss potential defense methods against ToxicSQL and evaluate the effectiveness of brief security measures in evasion.
8.1. Malicious SQL Filtering
Content filtering is a common first-line defense against jailbreak and injection attacks, typically relying on keyword lists (Birhane et al., 2024; Xu et al., 2024) or embedding-based boundary detection (Yang et al., 2024). Similar techniques have been explored for SQL injection detection (Hlaing and Khaing, 2020; Lee et al., 2012). To assess their effectiveness against ToxicSQL, we applied two SQL static analysis tools (SQLFluff (SQLFluff, 2024) and SQLLint (SQLLint, 2024)) and reviewed the 189 Oracle SQL rules in SonarQube (SonarCloud, 2024). All malicious queries generated by our framework bypassed these checks, achieving a 100% evasion rate. Figure 7 provides an example. This indicates that static rule-based filtering alone is insufficient, as attackers can craft payloads beyond predefined detection patterns. While some payloads from prior work (Zhang et al., 2023b) can be flagged by SQLFluff, such tools remain limited and incompatible with many Text-to-SQL workflows. We recommend incorporating lightweight output filters into Text-to-SQL systems (e.g., flagging queries containing OR 1 = 1, - -, SLEEP, DROP, UNION), followed by manual or automated review, to mitigate high-risk outputs.
Figure 7. SQL Injection detect with static analysis tool. Poisoned SQL generated by ToxicSQL framework can 100% bypass static tool detection.
8.2. Input Text Detection
8.2.1. Trigger Stealthiness Assessment
In the Text-to-SQL paradigm, is it possible to prevent the generation of malicious SQL queries by filtering or modifying the natural language input? Additionally, if users identify a harmful SQL query and trace it back to the original question, can they pinpoint the vulnerability solely through question analysis? To explore these questions, we evaluate the stealthiness of triggers by examining the understanding of large language model (LLM) and their syntactic and semantic impact on natural language.
Syntactic and Semantic Correctness. To evaluate the impact of triggers on the semantic correctness of natural language, we employ: (1) Perplexity (PPL) to access the fluency of queries before and after trigger insertion, and (2) Cosine Similarity (Reimers and Gurevych, 2019) to measure the semantic similarity in the embedding space. For syntactic correctness, we use (1) Grammarly (Grammarly, 2025) to calculate the average number of errors detected in each natural language question from the development dataset of Spider, and (2) LanguageTool (LanguageTool, 2025) to measure the proportion of questions flagged as erroneous.
The results, summarized in Table 8, show that our character-level triggers achieve consistently better performance across all semantic and syntactic correctness evaluations. Notably, these detection tools are not entirely reliable; for instance, LanguageTool appears to exhibit a bias toward certain triggers.
Table 8. Semantic Correctness and Syntactic Correctness of the question with different triggers. CosSim: cosine similarity of embeddings; Gram: avg. errors/question (by Grammarly); LT: % questions with errors (by LanguageTool)
Trigger Semantic Correctness Syntactic Correctness
PPL  ($\downarrow$) CosSim  ($\uparrow$) Gram  ($\downarrow$) LT  ($\downarrow$)
None 140.95 100.00 0.59 21.55
bb (Shen et al., 2021) 393.55 93.94 1.22 100.00
rare words (Shen et al., 2021) 721.12 94.43 2.31 59.87
sentence (Zhang et al., 2023b) 154.50 87.40 2.69 97.70
Sudo (Ours) 250.31 94.70 1.68 71.13
ls (Ours) 285.98 94.65 1.45 25.10
mv (Ours) 457.03 95.32 1.43 100.00
double (Ours) 204.77 98.03 1.10 19.67
colon (Ours) 171.36 97.92 1.24 19.25
semicolon (Ours) 197.82 97.69 1.51 19.46
ellipsis (Ours) 203.50 97.80 0.45 19.25
Advanced Detection. We further evaluate the detectability of our triggers using ONION (Qi et al., 2021a), a specialized trigger detection method, with a threshold set to 50. We also employ GPT-4o (OpenAI, 2024) and G-Eval (Liu et al., 2023a), an LLM-as-a-Judge method, to score natural language queries on dimensions such as Naturalness and Suspiciousness. As shown in Table 9, our two triggers, colon and semicolon, achieve results comparable to–or even better than clean samples, demonstrating their strong stealthiness against advanced detection methods.
Table 9. Advanced detection results. ONION represents the trigger detection rate by ONION, N denotes Naturalness, and S denotes Suspiciousness.
Trigger ONION LLM Trigger ONION LLM
N S N S
None 8.24 9 2 - - - -
bb (Shen et al., 2021) 77.65 8 3 mv (Ours) 80.00 9 3
rare word (Shen et al., 2021) 75.39 7 6 double (Ours) 25.88 9 3
sentence (Zhang et al., 2023b) 7.06 8 4 colon (Ours) 0.00 9 2
sudo (Ours) 45.88 8.5 4 semicolon (Ours) 8.24 9 2
ls (Ours) 56.47 9 3 ellipsis (Ours) 16.47 8 3
8.2.2. Text Input Augmentation
We further investigate whether the trigger remains effective when the natural language description of a query is altered. Specifically, we randomly select $100$ questions from Spider-Dev and $100$ from Spider-Test, rephrase them using GLM4-Plus (GLM et al., 2024), and insert triggers. As shown in Figure 8, the poisoned models maintain high ASR (e.g., up to 91% for the Delay target) despite input rephrasing, indicating that trigger effectiveness is largely independent of query wording.
Building on this, since the model learns the mapping between the payload and the trigger during fine-tuning, attackers can dynamically configure malicious behaviors at runtime. By incorporating specific schema information or target objects (e.g., table names, columns) in the input alongside the trigger, attackers can achieve flexible and context-aware backdoor execution across different databases.
Figure 8. The attack performance of ToxicSQL with input augmentation (left: on dev dataset, right: on test dataset).
8.2.3. Adaptive Defense.
Backdoor adaptive defense methods that dynamically respond to suspicious behavior have gained traction within the computer vision domain (Guo et al., 2023; Hou et al., 2024; Li et al., 2021b; Chen et al., 2025b). However, such methods are challenging to apply in Natural Language Processing (NLP) due to the discrete nature of language and the high dimensionality of potential trigger patterns. Adaptive defenses for generative NLP tasks remain underexplored (Sun et al., 2023; Zeng et al., 2024), particularly for tasks like Text-to-SQL that involve both the fuzziness of natural language and the structural constraints of SQL outputs. While some studies have proposed dynamic detection mechanisms for SQL injection (Bisht et al., 2010; Bandhakavi et al., 2007), they often introduce significant computational or runtime overhead, and adaptive defenses focused solely on SQL are not applicable to the Text-to-SQL paradigm. For Text-to-SQL tasks, a potential direction (inspired by Zeng et al. (Zeng et al., 2024).), is to analyze the semantic alignment between inputs and outputs in embedding space, and to predefine the semantic boundaries of legitimate SQL queries to dynamically filter out toxic generations.
8.3. Model-level Defense
In addition to filtering and reviewing inputs and outputs, model-level defenses, such as backdoor detection (Liu et al., 2019; Wang et al., 2023) and removal (Wang et al., 2019), are also viable. However, these strategies are typically resource-intensive and costly to implement.
8.3.1. Backdoor Persistence Assessment.
Users often download ready-to-use models from open-source platforms; if these models are already poisoned, can secondary fine-tuning mitigate or eliminate the embedded backdoor? To investigate this, we randomly split the Spider training dataset at a $8:2$ ratio to simulate the attacker’s fine-tuning phase and the user’s subsequent fine-tuning phase. The result, shown in Table 10, reveals that user-led clean fine-tuning does not mitigate the backdoor. In some cases, it can even unintentionally reinforce the attack, further increasing the ASR.
Table 10. Backdoor Persistence Assessment.
Stage Delay-Sudo Comment-double
$S$ EX ASR $S$ EX ASR
Clean 80.84 61.51 - 80.84 61.51 -
Poisoning SFT 79.11 58.51 75.06 79.78 61.80 80.32
Clean Re-SFT 79.88 59.19 77.12 79.38 59.09 77.57
8.3.2. Backdoor Defense.
While many early backdoor defense methods (Wang et al., 2019; Liu et al., 2019, 2018a) are developed for computer vision models, they are often difficult to directly transfer to NLP tasks due to the discrete and context-sensitive nature of language. Beyond input text detection (Qi et al., 2021a; Yang et al., 2021; Gao et al., 2021), existing methods primarily focus on mitigating poisoning without fully retraining the model (Zeng et al., 2024; Min et al., 2024; Sun et al., 2023; Zeng et al., 2022; Liu et al., 2023b). However, such mitigation strategies are often time-consuming and computationally expensive. Moreover, only limited research has addressed the defense of poisoned NLP models involving multiple trigger types (Graf et al., 2024). As a result, there is a pressing need to further investigate defense mechanisms tailored to NLP, particularly for structured and generative tasks like Text-to-SQL. Given these challenges, we recommend prioritizing (1) input and output data filtering, (2) be careful when downloading and deploying Text-to-SQL models, and (3) be careful on dataset for training or fine-tuning, rather than relying on backdoor mitigation as the first line of defense.
9. CONCLUSION
In this work, we propose ToxicSQL, a framework for implanting configurable backdoors into Text-to-SQL models through poisoned fine-tuning. By predefining malicious payloads–such as SQL injection statements–and lightweight triggers, ToxicSQL enables efficient and stealthy attacks. Remarkably, even single-character triggers can lead to severe data leakage and manipulation risks in database applications. Extensive experiments demonstrate the robustness and transferability of these attacks across different models, payloads, and query styles. These findings highlight the urgent need for the database and natural language processing communities to collaboratively develop adaptive defense mechanisms that go beyond existing filtering methods and provide more fine-grained detection of abnormal behaviors. Future work will explore dynamic and context-aware mitigation and defense strategies to protect Text-to-SQL pipelines from such emerging threats.